# LRT-Diffusion: Calibrated Risk-Aware Guidance for Diffusion Policies

### v1

@@ 5.3 Distributional Control and Return Bounds
$\varepsilon_{\rm in}:=\sup_{a\in\mathcal{S}(s)}|\hat{Q}(s,a)-Q^{\text{true}}(s,a)|,\quad\varepsilon_{\rm out}:=\sup_{a\notin\mathcal{S}(s)}|\hat{Q}(s,a)-Q^{\text{true}}(s,a)|,$
@@
Proposition 3 (Return comparison under offline errors).
@@
Decompose expectations into on-support and OOD parts. On-support deviations are bounded by $\varepsilon_{\rm in}$, off-support by $\varepsilon_{\rm out}$. Refer to Appendix  C.1 for complete proof. ∎
@@
Proposition 4 (From level-$\alpha$ to an OOD bound).
Under the assumption above, if each step’s gate has size at most $\alpha$ under $H_{0}$ and the sampler uses $T$ gated steps, then
$\eta(\pi_{\rm LRT})~\leq~\Pr(A)~\leq~1-(1-\alpha)^{T}~\approx~\alpha T\quad(\alpha T\ll 1).$
A is when at least one gate opens under $H_{0}$.
Remarks. The bound follows from a union/Bonferroni argument (or Ville’s inequality Ville1939 ) for the activation events. It formalizes the role of $\alpha$ as a risk knob: smaller $\alpha$ $\Rightarrow$ fewer/softer activations $\Rightarrow$ lower OOD rate, which—via Prop. 3—tightens a lower bound on the true return relative to pure $Q$-guidance when $\varepsilon_{\rm out}\!\geq\!\varepsilon_{\rm in}$.
5.4 A sufficient condition for LRT to dominate Q
Combining Prop. 3 and Prop. 4 yields
$\Delta_{Q^{\rm true}}\ \geq\ \Delta_{\hat{Q}}\ -\ 2\varepsilon_{\rm in}\ -\ \nu\big(\alpha T+\eta(\pi_{Q})\big).$
@@ 5.4 A sufficient condition for LRT to dominate Q
$\boxed{\alpha\ \leq\ \alpha_{\max}\ :=\ \frac{\Delta_{\hat{Q}}\ -\ 2\varepsilon_{\rm in}\ -\ \nu\,\eta(\pi_{Q})}{\nu\,T}\,,}$
@@
The bound is sufficient (and conservative): it uses the union proxy $\eta(\pi_{\rm LRT})\!\leq\!\alpha T$ and the critic gap $\Delta_{\hat{Q}}$, which may be biased off–support. Nevertheless it yields actionable levers: (i) decreasing $T$ (late–step gating) or $\beta_{\max}$ reduces $\nu$ and enlarges the feasible range; (ii) making $\eta(\pi_{Q})$ small (a conservative anchor for the QQ–step) increases $\alpha_{\max}$; (iii) evidence–tied guidance near $\mu_{\text{LRT}}$ can raise $\Delta_{\hat{Q}}$. In the favorable regime where $\eta(\pi_{Q})$ is small, $\alpha_{\max}$ grows as $\alpha_{\max}=(\Delta_{\hat{Q}}-2\varepsilon_{\rm in}-\nu\,\eta(\pi_{Q}))/(\nu T)$, so moderate $\alpha$ can still be certified. In practice we treat $\widehat{\alpha}_{\max}$ as a diagnostic and sweep $\alpha$ on a log grid; the selected $\alpha$ is the knee of the return–risk curve (Fig. 1), while the certificate provides a sanity upper bound. For further discussion, see Appendix C.5
6 Experiments
We empirically study whether LRT guidance delivers a calibrated, interpretable risk knob at inference and improves the return–OOD trade‑off over standard QQ‑guided sampling. Unless otherwise noted, diffusion training is vanilla (Sec. 4.2); all risk control is applied only at inference via the calibrated LRT gate.
6.1 Tasks and Datasets
@@
We report three metrics. (i) Return: mean $\pm$ s.e. over $N_{\mathrm{roll}}$ episodes; (ii) Realized Type‑I: under $H_{0}$ with the deployed sampler, the frequency $\Pr[\ell_{\mathrm{cum}}\geq\hat{\tau}$] (target $\alpha$), shown in Figure 1; (iii) State‑conditional OOD: a kk‑NN proxy on $(s,a)$: for each generated pair, take the $k$ nearest in state and flag OOD if the action distance exceeds the qq‑th percentile of in‑dataset radii (defaults k=50k{=}50, $q{=}95\%$).
@@ 6.2 Baselines
  • LRT+Q: LRT mean + small $Q$ step (calibration matches deployment).
6.3 Implementation & Reproducibility
Setup.
IQL critic ($\times2$56 MLP, $\gamma{=}0.99$, expectile 0.7); advantages on standardized $(s,a)$; labels: global top-$p$ (p=0.2p{=}0.2). Diffusion: T=50T{=}50, DDPM linear noise $(1{-}\alpha_{t})$ with endpoints $10^{-4}\!\to\!2{\times}10^{-2}$, MLP backbone (SiLU), two $\epsilon$ heads; AdamW $2{\times}10^{-4}$, batch 1024, 150 epochs. Class balancing via EMA positive rate; optional within-positive soft weights ($\tau_{A}$, $u_{\max}$). Full configs in Appx. E.
Inference & calibration.
Use reverse/posterior variance $\tilde{\sigma}_{t}^{2}$ for both heads and LLR; gate defaults $\beta_{\max}{=}1$, $\delta{\in}[1,2]$. Calibrate a single threshold $\hat{\tau}$ on a held-out state set by the fixed-point update
@@ 6.3 Implementation & Reproducibility
Each configuration: $5$ seeds $\times$ 1010 episodes; report mean $\pm$ s.e.; actions mapped back to env scale and clipped; using D4RL-raw returns. Calibration uses $n\in[3000,5000]$ reverse trajectories.
6.4 Main Results
We sweep $\alpha\!\in\!\{0.20,0.10,0.05,0.02,0.01\}$ with fixed $(\beta_{\max},\delta)$ and compare QG, LRT, and LRT+Q. LRT is expected to (i) honor the target $\alpha$ (realized Type‑I $\approx\alpha$ within DKW tolerance), and (ii) trade return for lower state‑conditional OOD as $\alpha$ decreases, yielding improved Pareto fronts relative to QG in regimes where the critic is unreliable off‑support.
@@ 6.4 Main Results
Table 1: Main results on D4RL MuJoCo. For each task/dataset and method, we report return (D4RL raw return) and state‑conditional OOD (%, lower is better). LRT honors the target $\alpha$ while improving the return–OOD trade‑off over QG in our implementation.
Data Name Mode Return (raw.) OOD
hopper-medium-replay-v2 lrt 377 0.017
hopper-medium-replay-v2 lrt+q 373 0.066
hopper-medium-replay-v2 q 366 0.023
halfcheetah-medium-replay-v2 lrt 578 0.011
halfcheetah-medium-replay-v2 lrt+q 786 0.094
halfcheetah-medium-replay-v2 q 706 0.092
walker2d-medium-replay-v2 lrt 362 0.003
walker2d-medium-replay-v2 lrt+q 436 0.033
walker2d-medium-replay-v2 q 433 0.032
Hopper
HalfCheetah
Walker2d
@@
Figure 1: Risk–performance and Pareto fronts across tasks. Left of each row: risk–performance curves versus target $\alpha$. We report return (top), realized Type-I (middle), and state-conditional OOD (bottom) for LRT (solid) and QG (dashed); the realized Type-I tracks the target (gray) within finite-sample DKW bands. Right of each row: Pareto fronts (OOD vs. return; color encodes $\alpha$). LRT shifts the frontier up-and-left relative to QG on tasks where off-support critic error dominates, yielding higher return at lower OOD for the same $\alpha$. Error bars denote standard errors over evaluation rollouts.
Results at a glance.
Table 1 and Fig. 1 tell a consistent story. Across tasks, LRT achieves the lowest state-conditional OOD while keeping return competitive; adding a small value step (LRT+Q) raises return but at a higher OOD cost. Concretely, on Hopper LRT (377, 1.7%) outperforms QG (366, 2.3%) and avoids the OOD increase of LRT+Q (373, 6.6%); on HalfCheetah LRT+Q attains the top return (786, 9.4%) while LRT is conservative (578, 1.1%) and QG sits between (706, 9.2%); on Walker2d LRT minimizes OOD (0.3%), LRT+Q maximizes return (436, 3.3%), and QG is similar (433, 3.2%). Thus, LRT provides a low-risk anchor and the $Q$-composition moves the policy along the return–risk frontier.
Reading the curves and fronts.
@@
In the left panels of Fig. 1, the realized Type-I (middle row) closely tracks the target $\alpha$ within DKW bands, validating calibration. As $\alpha$ decreases, return (top) drops modestly while state-conditional OOD (bottom) decreases consistently, exhibiting a smooth, monotone risk knob. The right panels (Pareto fronts) make the trade-off explicit: on Hopper and Walker2d, LRT shifts the frontier up-and-left relative to QG—higher return at lower OOD for a fixed $\alpha$—whereas on HalfCheetah, adding a small $Q$-step can push return further at the cost of OOD, matching the intended “anchor-plus-exploitation” behavior.
6.5 Ablations
Effect of $\alpha$.
Smaller $\alpha$ $\Rightarrow$ larger $\hat{\tau}$, fewer gate activations, lower state-conditional OOD with modest return drop (consistent with the left panels in Fig. 1).
Cap $\beta_{\max}$ & temperature $\delta$.
Decreasing $\beta_{\max}$ contracts updates toward $\mu_{u}$ (uniform OOD reduction even at large $\alpha$); decreasing $\delta$ sharpens the switch and approaches hard LRT (shifting Pareto up–left).
Labeling $p$.
Moderate $p\in[0.1,0.3]$ is robust; calibration preserves Type-I semantics regardless of $p$, so Pareto trends remain unchanged.
Evidence dynamics.
Prefix LLRs typically hover near zero at early high-noise steps and exceed $\hat{\tau}$ only later, if at all; the gate $\beta_{t}/\beta_{\max}$ remains near zero before crossing and then rises smoothly (Fig. 2). This matches the monotone risk knob seen in Fig. 1.
@@ 6.5 Ablations
Figure 2: LLR traces. Prefix LLR across denoising steps on random states; lines are different trajectories. Data: hopper-medium-replay-v2.
7 Conclusion and Future Work
@@
We presented LRT-Diffusion, a calibrated, inference-only guidance rule that turns diffusion-policy denoising into an evidence-gated process governed by a single, interpretable risk knob $\alpha$. Under equal covariances, the hard likelihood-ratio test (LRT) is uniformly most powerful at level $\alpha$; our smooth gate retains the same statistical semantics through a simple Monte-Carlo calibration matched to the deployed sampler. Empirically, LRT delivers calibrated control of Type-I risk and consistently improves the return–OOD trade-off over standard $Q$-guided sampling, while integrating seamlessly with a small critic step for additional exploitation. Together, these results demonstrate that risk-aware diffusion guidance can be achieved entirely at inference without modifying training objectives.
Several extensions naturally follow. (1) Learned or adaptive gating. Instead of a fixed calibrated threshold, the gate could be learned or amortized across tasks, or adapt a target $\alpha$ online under deployment constraints while preserving calibration. (2) Richer statistical models. Relaxing the equal-covariance assumption with heteroscedastic or learned head variances may yield stronger likelihood ratios and sharper decisions. (3) Sequential and early-stopping tests. Connecting the cumulative LLR to sequential or martingale tests (e.g., SPRT Wald1945SPRT ) would enable any-time risk control and early termination of denoising when the evidence is decisive, reducing compute without violating the level-$\alpha$ guarantee. (4) Broader integration and evaluation. Composing LRT with model-based planners or diffusion world models could link evidence-gated actions to trajectory-level reasoning, potentially higher returns on broader datasets.
Appendix A Gate Selection
A.1 Hard Gate-UMP: Proof of Proposition 1
@@
Let $\Phi$ be the class of (possibly randomized) tests $\phi:\mathcal{A}^{T+1}\!\to[0,1]$ with size at most $\alpha$, i.e. $\mathbb{E}_{H_{0}}[\phi]\leq\alpha$. By the Neyman–Pearson lemma for simple $H_{0}$ vs. $H_{1}$, there exists $k>0$ such that the likelihood-ratio test $\phi^{\star}(a_{0:T})=\mathbf{1}\{L(a_{0:T})\geq k\}$ (with possible randomization on the boundary $\{L=k\}$) attains size exactly $\alpha$ and maximizes the power $\mathbb{E}_{H_{1}}[\phi]$ over $\Phi$. Since $\log(\cdot)$ is monotone, thresholding $L$ is equivalent to thresholding $\ell_{\mathrm{cum}}=\log L$. Hence the stated test is most powerful at level $\alpha$.
@@
A.2 Soft to Hard Limit: Proof of Lemma  1
@@
$\beta_{t}(\ell_{\mathrm{cum}};\tau,\delta,\beta_{\max})~=~\beta_{\max}\,\sigma\!\Big(\tfrac{\ell_{\mathrm{cum}}-\tau}{\delta}\Big),\qquad\sigma(x)=\tfrac{1}{1+e^{-x}},$
@@
$\lim_{\delta\downarrow 0}\ \sigma\!\Big(\tfrac{x-\tau}{\delta}\Big)~=~\mathbf{1}\{x\geq\tau\}.$
@@
$\beta_{t}(\ell_{\mathrm{cum}};\tau,\delta,\beta_{\max})~\xrightarrow[\delta\downarrow 0]{}~\beta_{\max}\,\mathbf{1}\{\ell_{\mathrm{cum}}\geq\tau\}.$
@@
$\mu_{t}^{\mathrm{soft}}~=~\mu_{u}+\beta_{t}(\mu_{c}-\mu_{u})~\xrightarrow[\delta\downarrow 0]{}~\mu_{u}+\beta_{\max}\,\mathbf{1}\{\ell_{\mathrm{cum}}\geq\tau\}\,(\mu_{c}-\mu_{u})~=:\ \mu_{t}^{\mathrm{hard}}.$
@@
$\mu_{t}^{\mathrm{soft},(\delta)}~\longrightarrow~\mu_{t}^{\mathrm{hard}}\quad\text{a.s.}$
@@
$a_{t-1}^{\mathrm{soft}}=\mu_{t}^{\mathrm{soft}}+\sigma_{t}z_{t}~\longrightarrow~\mu_{t}^{\mathrm{hard}}+\sigma_{t}z_{t}=a_{t-1}^{\mathrm{hard}}\quad\text{a.s.}$
@@
Appendix B Proofs for Finite-sample Calibration and Stability

### v2

@@ 5.3. Distributional Control and Return Bounds
$\tiny\varepsilon_{\rm in}:=\sup_{a\in\mathcal{S}(s)}|\hat{Q}(s,a)-Q^{\text{true}}(s,a)|,\quad\varepsilon_{\rm out}:=\sup_{a\notin\mathcal{S}(s)}|\hat{Q}(s,a)-Q^{\text{true}}(s,a)|,$
@@
[…]
@@
Decompose expectations into on-support and OOD parts. On-support deviations are bounded by $\varepsilon_{\rm in}$, off-support by $\varepsilon_{\rm out}$. Refer to Appendix D.1 for complete proof. ∎
@@
[…]
Assuming monotone support, let $\pi_{\rm LRT}$ be the LRT-gated policy with cumulative LLR threshold $\tau$ calibrated under $H_{0}$ at level $\alpha$. Then the expected state-conditional OOD rate satisfies
$\eta(\pi_{\rm LRT})\;=\;\mathbb{E}_{s}\big[\Pr_{a_{0}\sim\pi_{\rm LRT}(\cdot|s)}[a_{0}\notin\mathcal{S}(s)]\big]\;\lesssim\;\Pr_{H_{0}}(\ell_{\rm cum}\geq\tau)\;\approx\;\alpha.$
Here, $\ell_{\rm cum}$ is the cumulative log-likelihood ratio over the reverse trajectory, and the approximation holds up to finite-sample calibration error (Thm. 5.2.1). Proof see Appendix D.2
Thus, smaller $\alpha$ directly yields fewer activations and lower expected OOD.
5.4. A sufficient condition for LRT to dominate Q
Combining Prop. 5.3 and Prop. 5.3 yields
$\Delta_{Q^{\rm true}}\ \geq\ \Delta_{\hat{Q}}\ -\ 2\varepsilon_{\rm in}\ -\ \nu\big(\alpha+\eta(\pi_{Q})\big).$
@@ 5.4. A sufficient condition for LRT to dominate Q
$\alpha\ \leq\ \alpha_{\max}\ :=\ \frac{\Delta_{\hat{Q}}\ -\ 2\varepsilon_{\rm in}\ -\ \nu\,\eta(\pi_{Q})}{\nu\,}\,$,
@@
The bound is sufficient (and conservative): it uses the calibrated bound $\eta(\pi_{\rm LRT})\!\lesssim\!\alpha$ and the critic gap $\Delta_{\hat{Q}}$, which may be biased off–support. Nevertheless it yields actionable levers: (i) decreasing $\beta_{\max}$ reduces $\nu$ and enlarges the feasible range; (ii) making $\eta(\pi_{Q})$ small (a conservative anchor for the QQ–step) increases $\alpha_{\max}$; (iii) evidence–tied guidance near $\mu_{\text{LRT}}$ can raise $\Delta_{\hat{Q}}$. In the favorable regime where $\eta(\pi_{Q})$ is small, $\alpha_{\max}$ grows as $\alpha_{\max}=(\Delta_{\hat{Q}}-2\varepsilon_{\rm in}-\nu\,\eta(\pi_{Q}))/\nu$, so moderate $\alpha$ can still be certified. In practice we treat $\widehat{\alpha}_{\max}$ as a diagnostic and sweep $\alpha$ on a log grid; the selected $\alpha$ is the knee of the return–risk curve (Fig. 2), while the certificate provides a sanity upper bound. For further discussion, see Appendix D.6
6. Experiments
We empirically study whether LRT guidance delivers a calibrated, interpretable risk knob at inference time and yields a calibrated return–risk frontier relative to standard $Q$-guided sampling. Unless otherwise noted, diffusion training is vanilla (Sec. 4.2); all risk control is applied only at inference via the calibrated LRT gate.
6.1. Tasks and Datasets
@@
[…]
@@ 6.2. Baselines
  • LRT+Q: LRT mean + small $Q$ step.
6.3. Implementation & Reproducibility
Setup. IQL critic ($\times2$56 MLP, $\gamma{=}0.99$, expectile 0.7); advantages on standardized $(s,a)$; labels: global top-$p$ (p=0.2p{=}0.2). Diffusion: T=50T{=}50, DDPM linear noise $(1{-}\alpha_{t})$ with endpoints $10^{-4}\!\to\!2{\times}10^{-2}$, MLP backbone (SiLU), two $\epsilon$ heads; AdamW $2{\times}10^{-4}$, batch 1024, 150 epochs. Class balancing via EMA positive rate; optional within-positive soft weights ($\tau_{A}$, $u_{\max}$). Full configs in Appx. F.
Inference & calibration. Use reverse/posterior variance $\tilde{\sigma}_{t}^{2}$ for both heads and LLR; gate defaults $\beta_{\max}{=}1$, $\delta{\in}[1,2]$. Calibrate a single threshold $\hat{\tau}$ on a held-out state set by the fixed-point update
@@ 6.3. Implementation & Reproducibility
Protocol & compute. Each configuration: $10$ seeds $\times$ $10$ episodes over the entire training and evaluation time; report mean $\pm$ std; actions mapped back to env scale and clipped; using D4RL-raw returns. Calibration uses $n\in[3000,5000]$ reverse trajectories.
6.4. Main Results
We perform structured hyperparameter sweeps to characterize the return–risk frontier. For LRT, we sweep the risk level $\alpha\in\{0.20,0.10,0.01,0.005,0.001\}$ with fixed $(\beta_{\max},\delta)$. For QG, we sweep the maximum guidance step $\lambda_{\max}\in\{0.2,0.1,0.05,0.02,0.005\}$ (same schedule and clipping). For LRT+Q, we sweep the Cartesian product $\alpha\times\lambda_{\max}$ (25 settings per seed).
[…]
@@ 6.4. Main Results
Table 1. Main results on D4RL MuJoCo. For each task/dataset and method, we report return (D4RL raw return) and state-conditional OOD (reported in units of $\times 10^{-2}$; lower is better). Results correspond to the operating point that maximizes mean return across seeds within each method’s sweep.
[…]
hopper-medium-replay-v2
Refer to caption Refer to caption
halfcheetah-medium-replay-v2
Refer to caption Refer to caption
walker2d-medium-replay-v2
Refer to caption Refer to caption
@@
Figure 2. Risk–performance and Pareto fronts across tasks. Left of each row: risk–performance curves versus target $\alpha$. We report return (top), realized Type-I (middle), and state-conditional OOD (bottom) for LRT (solid) and QG (dashed); the realized Type-I tracks the target (gray) within finite-sample DKW bands.Right of each row: Pareto fronts (OOD vs. return; color encodes $\alpha$). LRT shifts the frontier up-and-left relative to QG on tasks where off-support critic error dominates, yielding higher return at lower OOD for the same $\alpha$. Error bars denote standard errors over evaluation rollouts.
Results at a glance. Across tasks, LRT often yields the lowest state-conditional OOD, acting as a conservative, low-risk anchor. Adding a small critic step (LRT+Q) typically increases return but can increase OOD, tracing a return–risk frontier. Notably, on medium-replay datasets LRT’s OOD advantage is most pronounced, while on medium datasets value guidance often delivers larger return gains. See Appx. G for per-task interpretation.
@@
Reading the curves and fronts. In the left panels of Fig. 2, the realized Type-I (middle row) closely tracks the target $\alpha$ within DKW bands, validating calibration. As $\alpha$ decreases, return (top) drops modestly while state-conditional OOD (bottom) decreases consistently, exhibiting a smooth, monotone risk knob. The right panels (Pareto fronts) make the trade-off explicit: on Hopper and Walker2d, LRT shifts the frontier up-and-left relative to QG—higher return at lower OOD for a fixed $\alpha$—whereas on HalfCheetah, adding a small $Q$-step can push return further at the cost of OOD, matching the intended “anchor-plus-exploitation” behavior.
6.5. Ablations
Effect of $\alpha$: Smaller $\alpha$ $\Rightarrow$ larger $\hat{\tau}$, fewer gate activations, lower state-conditional OOD with modest return drop (consistent with the left panels in Fig. 2). Cap $\beta_{\max}$ & temperature $\delta$: Decreasing $\beta_{\max}$ contracts updates toward $\mu_{u}$ (uniform OOD reduction even at large $\alpha$); decreasing $\delta$ sharpens the switch and approaches hard LRT (shifting Pareto up–left). Labeling $p$: Moderate $p\in[0.1,0.3]$ is robust; calibration preserves Type-I semantics regardless of $p$, so Pareto trends remain unchanged (see Appendix H.1). Evidence dynamics: Prefix LLRs typically hover near zero at early high-noise steps and exceed $\hat{\tau}$ only later, if at all; the gate $\beta_{t}/\beta_{\max}$ remains near zero before crossing and then rises smoothly (Fig. 3). This matches the monotone risk knob seen in Fig. 2. Stress test under variance mismatch: We conduct a stress test under deliberate variance mis-specification of the background head at inference. Increasing mismatch inflates Type-I error and OOD, causing return degradation; matched variance ($s=1$) preserves risk control, while severe mismatch ($s\geq 2$) degrades guarantees. See Appendix H.2 for details.
@@ 6.5. Ablations
Figure 3. LLR traces. Prefix LLR across denoising steps on random states; lines are different trajectories. Data: hopper-medium-replay-v2.
7. Conclusion
@@
We presented LRT-Diffusion, a calibrated, inference-only guidance rule that turns diffusion-policy denoising into an evidence-gated process governed by a single, interpretable risk knob $\alpha$. Under equal covariances, the hard likelihood-ratio test (LRT) is uniformly most powerful at level $\alpha$; our smooth gate retains the same statistical semantics through a simple Monte-Carlo calibration matched to the deployed sampler. Empirically, LRT delivers calibrated control of Type-I risk and often improves the return–OOD trade-off over standard $Q$-guided sampling; moreover, in the experiments in Table 1, the composed sampler (LRT+Q) consistently achieves the highest return, highlighting that evidence-gated guidance can stabilize and amplify the benefits of value-gradient updates. Together, these demonstrate that risk-aware diffusion guidance can be achieved entirely at inference without modifying training objectives.
Appendix A Detailed pipeline of LRT-Diffusion
For a detailed illustration of the inference-time pipeline, see Fig. 4 in the Appendix.
Figure 4. Overall pipeline of LRT-Diffusion (reproduced for reference). Training is vanilla; all risk control is applied at inference via a calibrated likelihood-ratio gate.
Appendix B Gate Selection
B.1. Hard Gate-UMP: Proof of Proposition 5.1
@@ Appendix B Gate Selection
Let $\Phi$ be the class of (possibly randomized) tests $\phi:\mathcal{A}^{T+1}\!\to[0,1]$ with size at most $\alpha$, i.e. $\mathbb{E}_{H_{0}}[\phi]\leq\alpha$. By the Neyman–Pearson lemma for simple $H_{0}$ vs. $H_{1}$, there exists $k>0$ such that the likelihood-ratio test $\phi^{\star}(a_{0:T})=\mathbf{1}\{L(a_{0:T})\geq k\}$ (with possible randomization on the boundary $\{L=k\}$) attains size exactly $\alpha$ and maximizes the power $\mathbb{E}_{H_{1}}[\phi]$ over $\Phi$. Since $\log(\cdot)$ is monotone, thresholding $L$ is equivalent to thresholding $\ell_{\mathrm{cum}}=\log L$. Hence the stated test is most powerful at level $\alpha$. ∎
@@
B.2. Soft to Hard Limit: Proof of Lemma 5.1
@@
$\beta_{t}(\ell_{\mathrm{cum}};\tau,\delta,\beta_{\max})\penalty 10000\ =\penalty 10000\ \beta_{\max}\,\text{Sigmoid}\!\left(\tfrac{\ell_{\mathrm{cum}}-\tau}{\delta}\right),\qquad\text{Sigmoid}\!\left(x\right)=\tfrac{1}{1+e^{-x}},$
@@
$\lim_{\delta\downarrow 0}\ \text{Sigmoid}\!\left(\tfrac{x-\tau}{\delta}\right)\penalty 10000\ =\penalty 10000\ \mathbf{1}\{x\geq\tau\}.$
@@
$\beta_{t}(\ell_{\mathrm{cum}};\tau,\delta,\beta_{\max})\penalty 10000\ \xrightarrow[\delta\downarrow 0]{}\penalty 10000\ \beta_{\max}\,\mathbf{1}\{\ell_{\mathrm{cum}}\geq\tau\}.$
@@
$\mu_{t}^{\mathrm{soft}}\penalty 10000\ =\penalty 10000\ \mu_{u}+\beta_{t}(\mu_{c}-\mu_{u})\penalty 10000\ \xrightarrow[\delta\downarrow 0]{}\penalty 10000\ \mu_{u}+\beta_{\max}\,\mathbf{1}\{\ell_{\mathrm{cum}}\geq\tau\}\,(\mu_{c}-\mu_{u})\penalty 10000\ =:\ \mu_{t}^{\mathrm{hard}}.$
@@
$\mu_{t}^{\mathrm{soft},(\delta)}\penalty 10000\ \longrightarrow\penalty 10000\ \mu_{t}^{\mathrm{hard}}\quad\text{a.s.}$
@@
$a_{t-1}^{\mathrm{soft}}=\mu_{t}^{\mathrm{soft}}+\sigma_{t}z_{t}\penalty 10000\ \longrightarrow\penalty 10000\ \mu_{t}^{\mathrm{hard}}+\sigma_{t}z_{t}=a_{t-1}^{\mathrm{hard}}\quad\text{a.s.}$
@@
Appendix C Proofs for Finite-sample Calibration and Stability